Required Informatlon The following Information applies to the questions displayed below Three different companies each purchased trucks on January 1, Year 1, for $86,000. Each truck was expected to last four Years or 250,000 mlles. Salvage value was estimated to be $4,000. All three trucks were driven 76,000 miles In Year 1, 56,000 miles In Year 2, 51,000 miles In Year 3, and 71,000 miles In Year 4. Each of the three companles earned $75,000 of cash revenue during each of the four Years. Company A uses straight-lIne depreclation, company B uses double- declining-balance depreciation, and company C uses units-of-production depreciation. Answer each of the following questions. Ignore the effects of Income taxes. b-1. Calculate the net Income for Year 4 b-2. Which company will report the lowest amount of net Income for Year 4? Complete this question by entering your answers in the tabs below. Required B1 Required B2 Calculate the net income for Year 4. (Round your Per Unit Cost to 3 decimal places.) Net Income Company A Company B Company C < Required B1 Required B2>
Answer #1

b-1.

Net Income
Company A $       54,500
Company B $       68,250
Company C $       52,067

b-2. Company C will report the lowest amount of net income for Year 4.

Working:

Company A:

Depreciation = ($86000 - $4000)/4 = $82000/4 = $20500

Net income = $75000 - $20500 = $54500

Company B:

Straight-line depreciation rate = 100%/4 = 25%

Double declining balance depreciation rate = 2 x 25% = 50%

Depreciation for year 1 = 50% x $86000 = $43000

Depreciation for year 2 = 50% x ($86000 - $43000) = 50% x $43000 = $21500

Depreciation for year 3 = 50% x ($43000 - $21500) = 50% x $21500 = $10750

Depreciation for year 4 = ($21500 - $10750) - $4000 = $6750

Net income = $75000 - $6750 = $68250

Company C:

Depreciation per mile = ($86000 - $4000)/(76000 + 56000 + 51000 + 71000) = $82000/254000 = 0.3228 = $0.323 per mile

Depreciation = 71000 x $0.323 = $22933

Net income = $75000 - $22933 = $52067